OpenStudy (madmerc):

Simplify 4 over -2 - 2i -1 - i -2 - i -1 + i -2 + i

Nnesha (nnesha):

\[\huge\rm \frac{ 4 }{ -2-2i }\] multiply top and bottom of the fraction by the conjugate of the denominator

OpenStudy (madmerc):

what does that mean?

Nnesha (nnesha):

do you know what conjugate is ?

OpenStudy (madmerc):

no

Nnesha (nnesha):

conjugate of a +bi is `a-bi` change the sign of imaginary term

Nnesha (nnesha):

so what's the conjugate of `-2-2i` ?

OpenStudy (madmerc):

-2+2i?

Nnesha (nnesha):

right now multiply numerator and denominator by `-2+2i`\[\large\rm \frac{ 4 }{ -2-2i }*\frac{-2+2i}{-2+2i}\]

Nnesha (nnesha):

familiar with the distributive property ?

OpenStudy (theloshua):

lol dude basically you flip the middle sign of the denometer and multiply its by the numerator

Nnesha (nnesha):

multiply by both *denominator and numerator *

OpenStudy (madmerc):

I still dont understand.

Nnesha (nnesha):

it's like rationalize the denominator when we have to move radical sign from bottom to the top

Nnesha (nnesha):

to move the imaginary term from the denominator we should multiply top and bottom by the `conjugate ` of the denominator (which is -2-2i)

Nnesha (nnesha):

conjugate of -2-2i is `-2+2i` right so now multiply both top and bottom by -2+2i make sense ?

OpenStudy (madmerc):

so... 4 and -2+2i times -2+2i?

Nnesha (nnesha):

right \[\large\rm \frac{ 4 }{ -2-2i }*\frac{-2+2i}{-2+2i}\] which can be written as \[\large\rm \frac{ 4(-2+2i) }{ (-2-2i)(-2+2i) }\]

OpenStudy (madmerc):

then what?

Nnesha (nnesha):

now apply distributive property 4(-2+2i) and foil (-2-2i)(-2+2i)

OpenStudy (madmerc):

so -8+8i over 4 - 2i ?

Nnesha (nnesha):

distributive property \[\rm a(b+c)=a*b+a*c=ab+ac\]

Nnesha (nnesha):

-8+8i is correct but how did you get 4-2i ?

Nnesha (nnesha):

remember \[\large\rm i= \sqrt{-1} ~~~and ~~~~~~~~~i^2= -1\]

OpenStudy (madmerc):

I don't understand the bottom part I tried foil but I don't get it and I need to finish this test

Nnesha (nnesha):

how did you foil it. show your work . i'll try to find out the mistake

OpenStudy (madmerc):